OpenStudy (anonymous):

What is the y-intercept of the function? y = 20,000(0.95)x

OpenStudy (anonymous):

y int is when x=0

OpenStudy (anonymous):

the answer is 0 :)

OpenStudy (anonymous):

if u do not have a +something aka b value than it is always 0 :)

OpenStudy (anonymous):

no i don't think so

OpenStudy (anonymous):

\[y = 20,000(0.95)^x \] now let \(x=0\) get \[y = 20,000(0.95)x^0=20,000\] for an exponential function like the the y intercept is the number out front